A cheetah is hunting. Its prey runs for 3.3 s at a constant velocity of +9.7 m/s. Starting from rest, what constant acceleration must the cheetah maintain in order to run the same distance as its prey runs in the same time?
Answer #1

Here ,

let the acceleration of the cheetah is a

Using second equation of motion

0.50 * at^2 = u * t

0.50 * a * 3.3^2 = 3.3 * 9.7

solving for a

a = 5.88 m/s^2

the constant acceleration is 5.88 m/s^2